• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 21:25:06 on July30,2003

    Entry number 108223

    This entry is a followup to: 108222

    keyword=only 2 lasers were on for last NMR


    Looking at the strip tool for target temperatures, and then checking the laser
    screen, I realize that for the last NMR only Lasers 4,6 had been on. Laser 7 was "Enabled" but apparently I failed to click Start properly. I have started laser 7
    as/of 21:14.